Express −1+i in polar form. Hence show that (−1+i)16 is real and that 1/(−1+i)6 is purely imaginary, giving the value for each.
1
Expert's answer
2020-05-25T20:53:20-0400

"-1+i=\sqrt{2}(\cos \frac{3\pi}{4}+i\sin \frac{3\pi}{4})"

"(-1+i)^{16}=256(\cos \frac{48\pi}{4}+i\sin \frac{48\pi}{4})=256(\cos 12\pi+i\sin 12\pi)=256."

"\frac{1}{(-1+i)^6}=(-1+i)^{-6}=\frac{1}{8}(\cos \frac{18\pi}{4}+i\sin \frac{18\pi}{4})=\frac{1}{8}(\cos \frac{9\pi}{2}+i\sin \frac{9\pi}{2})=\frac{1}{8}i".